# The Key Physics of Ice Premelting

### v1

@@ I Introduction
On the 12th of march 1936 a team of Japanese researchers led by Ukichiro Nakaya succeeded in creating the first ever artificial snow crystals in a snow chamber under controlled conditions.Nakaya (1954) After several years of further research, the results were summarized and collected in the diagram that now bears his name.Nakaya (1954); Kobayashi (1957); Frank (1982); Furukawa and Wettlaufer (2007); Libbrecht (2022) The Nakaya diagram maps crystal habits as a function of temperature and humidity of the surrounding atmosphere (Figure 1). The complex structures that form as supersaturation is increased can be relatively well understood under the framework of crystal growth science.Frank (1982); Nelson (2001); Libbrecht (2022) However, the surprising transformation of solid hexagonal prisms from plate like to columnar like and back to plate like as temperature decreases has remained a mystery ever since.Frank (1982); Furukawa and Wettlaufer (2007); Libbrecht (2022)
@@
Kuroda went on to conduct ellipsometric and x-ray diffraction studies that were among the first few direct evidences for the presence of a quasi-liquid layer of melt water on the ice surface.Furukawa, Yamamoto, and Kuroda (1987); A. Kouchi and Kuroda (1987) Following these studies, the field experienced a surge of renewed activity, with numerous measurements and estimates of the quasi-liquid layer thickness carried out using a variety of modern spectroscopic and mechanical techniques.A. Kouchi and Kuroda (1987); Elbaum, Lipson, and Dash (1993); Lied, Dosch, and Bilgram (1994); Dosch, Lied, and Bilgram (1995, 1996); Wei, Miranda, and Shen (2001); Döppenschmidt and Butt (2000); Bluhm et al. (2002); Sadtchenko and Ewing (2002); Gelman Constantin et al. (2018); Mitsui and Aoki (2019)
@@
Interestingly, once the presence of liquid water droplets on the ice surface was confirmed by direct optical microscopy imaging, Faraday’s original hypothesis came under renewed scrutiny.Knight (1996) More recently, Sazaki and collaborators–at the Institute for Low Temperature Science, founded by Nakaya and later home to Kuroda–have raised serious concerns about this hypothesis, while studies on the ice surface using macroscopic probes have lead Bonn et al. to question the significance of ice premelting altogether.Liefferink et al. (2021); Demmenie et al. (2022, 2025)
@@
In this perspective article I discuss how a selected blend of concepts from these wide range of fields together with the help of computer simulations,Benet et al. (2016, 2019); MacDowell (2019); Llombart et al. (2019, 2020); Llombart, Noya, and MacDowell (2020); Sibley et al. (2021); Luengo-Márquez and MacDowell (2021); Luengo-Márquez, Izquierdo-Ruiz, and MacDowell (2022); Baran et al. (2022); Baran and MacDowell (2024); Baran et al. (2024); Baran, Llombart, and MacDowell (2025) allows to interpret and reconcile a large body of apparently conflicting empirical evidence into a unified and consistent picture of the structure and dynamics of premelting films.
Often, this description requires delving into rather intricate and subtle discussions. The readers less interested in the finer details of wetting physics and crystal growth can get an overall impression of the field by going through sections II, VI, IX, XI and XII only, while a read through section XII may be sufficient to get a broad picture of the current understanding and future challenges.
Figure 1: Diagram of snow crystal morphology, a descendant of the early Nakaya diagram due to Kobayashi.Kobayashi (1957) The figure displays the habit of ice crystals grown in the atmosphere as a function of temperature and excess water vapor density over ice saturation. The origin of the complex shapes that appear as saturation increases are relatively well understood. The reason why simple hexagonal prisms grown at low vapor saturation change from plates, to columns and back again to plates as temperature decreases has remained a mystery ever since Nakaya first published his results. Reproduced from Furukawa and Wettlaufer, Phys. Today 60 70 (2007),Furukawa and Wettlaufer (2007) with the permission of the American Institute of Physics.
@@ II Modern empirical evidence of ice premelting
The experimental measurement of ice premelting thickness has been reviewed in a large number of papers.Elbaum, Lipson, and Dash (1993); Petrenko (1994); Sadtchenko and Ewing (2002); Bartels-Rausch et al. (2014); Asakawa et al. (2016); Gelman Constantin et al. (2018); Slater and Michaelides (2019) All of them agree that the data exhibit enormous dispersion, with discrepancies spanning several orders of magnitude. However, the large extent of scatter can be reduced significantly by a judicious selection of experimental results. As noted in Ref.Dash, Rempel, and Wettlaufer (2006) there is no reason to expect that the premelting film thickness should be equal for surface and interfacial premelting. Indeed, computer simulations show that interfacial premelting structure is similar to that of surface premelted films only for very hydrophobic substrates.Baran, Llombart, and MacDowell (2025) Considering the case of surface premelting alone reduces signficantly the number of available measurements.Golecki and Jaccard (1978); Beaglehole and Nason (1980); Furukawa, Yamamoto, and Kuroda (1987); Elbaum, Lipson, and Dash (1993); Dosch, Lied, and Bilgram (1995); Döppenschmidt and Butt (2000); Sadtchenko and Ewing (2002); Gelman Constantin et al. (2018); Mitsui and Aoki (2019) However, out of all these studies, the early results by Golecki and Jacard using proton backscattering are clearly outliers,Golecki and Jaccard (1978) and can be ruled out. Furthermore, studies using Atomic Force Microscopy are necessarily invasive, and are prone to capillary condensation in between the tip and the surface,Pickering et al. (2018) so they are left out too. A final selection accounting for all the remaining measurements since 1990, using widely different non-invasive techniques such as ellipsometry, x-ray scattering, and infra-red, sum frequency or fluctuation spectroscopy lead to a rather different picture, with results differing by at most one order of magnitude for temperatures below about $1^{\circ}$C (Figure 2). The scatter remains bracketed within one order of magnitude even up to temperatures of $0.1^{\circ}$C if the results of Ref.Dosch, Lied, and Bilgram (1996) are discarded. All of these experimental measurements confirm that a premelting layer appears on the ice surface, with thicknesses that increase from the sub-nanometer scale to at most 10 nm in the temperature interval between 20-20 and $1^{\circ}$C. This appears somewhat higher but within reasonable agreement with computer simulation results for different water models, which consistently show the formation of a disordered liquid layer on the ice surface.Conde, Vega, and Patrykiejew (2008); Pfalzgraff, Neshyba, and Roeselova (2011); Limmer and Chandler (2014); Benet et al. (2016); Sánchez et al. (2017); Louden and Gezelter (2018); Qiu and Molinero (2018); Pickering et al. (2018); Kling, Kling, and Donadio (2018); Benet et al. (2019); Llombart et al. (2019, 2020); Llombart, Noya, and MacDowell (2020) Indeed, the extent of ice premelting can be quantified in computer simulations using suitable order parameters that allow to distinguish whether a water molecule is found in a liquid like or solid like environment (Figure 3).Lechner and Dellago (2008); Nguyen and Molinero (2015) Using such parameters, simulation studies report film thicknesses that lie in the sub-nanometer range up to about $1^{\circ}$C below the model’s triple point, and then grow somewhat above the nanometer thickness at higher temperatures.Conde, Vega, and Patrykiejew (2008); Limmer and Chandler (2014); Qiu and Molinero (2018); Pickering et al. (2018); Llombart et al. (2019, 2020); Llombart, Noya, and MacDowell (2020); Berrens, Bononi, and Donadio (2022)
@@ III A toy model of (complete) surface melting
Figure 4: Phase diagram of water in the neighborhood of the triple point. Full black lines illustrate the equilibrium phase boundaries, including the sublimation line separating ice from vapor and the condensation line, separating liquid water from vapor. The black dashed line is the metastable prolongation of the condensation line. Equilibrium premelting occurs for a path along the sublimation line, as indicated by the blue arrow. The equilibrium thickness of the premelting film is dictated by the distance of the sublimation line to the metastable prolongation of the condensation line, as indicated by the green segment. The red line indicates a non-equilibrium path, with three different regimes as described in Section IX. In region a, below the kinetic condensation line, shown in dashed blue, ice grows but the premelting film remains in steady state of constant thickness. In region b, above the kinetic condensation line but below the kinetic spinodal line shown in dashed red, the premelting film remains in steady state but vapor can condense and form droplets atop. In region c, above the kinetic spinodal line, condensation occurs faster than freezing, the premelting film diverges and ice freezes in a wet mode.
@@
Premelting, viewed as an equilibrium phenomenon, is the study of how the metastable liquid phase ($l$) intrudes at the interface between the coexisting bulk solid ($s$) and vapor phases ($v$) in the vicinity of a metastable liquid-vapor coexistence line (Fig. 3). For many substances, this can only occur as the system approaches the triple point along the sublimation line. For the special case of ice, however, notice the metastable prolongation of the liquid-vapor coexistence line runs very close to the sublimation line all the way down to the triple point (Fig. 4).
@@
Within the square gradient approximation, the excess surface free energy on the interface may be obtained in terms of $\phi(z)$, as a sum of bulk free energies adopted by the system along the interface. An additional square gradient term accounts for the cost of creating the inhomogeneity:
@@
A simple way to remedy this deficiency is to modify the biquadratic potential to account for the intrusion of a third phase with an order parameter in between those of the bulk solid and liquid phases, i.e. $f(\phi)=(1-\phi^{2})^{2}(t+\phi^{2})$, where $t\propto 1-T/T_{t}$ is a measure of the under-cooling (Figure 5-a). The Euler-Lagrange equation, Eq.(2) can be solved exactly for this potential, yielding:Lipowsky and Speth (1983)
@@
This result nicely describes the essence of premelting. For deep under-cooling, $t$ is much larger than $1-\tanh^{2}(\kappa z)$ for all $z$, and Eq.(4) becomes virtually equal to the sharp $\tanh(\kappa z)$ interfacial profile that is typical for two phases at coexistence. When the system approaches the triple point, however, $t\to 0$, the order parameter profile adopts a value $\phi(z)\approx\sqrt{t}$ close to zero as long as $1-\tanh^{2}(\kappa z)$ in the denominator remains larger than $t$. This occurs in an increasingly larger range of $z$ values as $t$ vanishes. Accordingly, the model predicts the gradual increase of the liquid phase in between the solid and the vapor phase on approaching the triple point (See Figure 5-b).
@@
Using the above results it is also possible to discuss what precisely can be understood as the onset of premelting.Qiu and Molinero (2018); Slater and Michaelides (2019); Nagata et al. (2019) The model shows that premelting appears in a continuous fashion, as a gradual flattening of the interfacial profile at the interface position $z=0$. i.e., there is no particular thermodynamic signature in the density profile of the solid-vapor interface signaling a premelting threshold. The best that can be done is to notice that the flattening of the interfacial density is the result of the stabilization of the liquid phase at a metastable minimum of the free energy surface. This is the only feature that can be identified as the theoretical onset of surface premelting. However, the signature of the metastable minimum is hardly detectable in the order parameter profile until the minima of $f(\phi)$ become sufficiently close in energy, i.e. $t\approx 0$. Therefore, for practical matters the onset cannot be identified until the change in slope and the emerging liquid film at $z=0$ is above some detection threshold of the experimental probe used in experiment.
Figure 5: Toy model of surface premelting. (a) Free energy of a two phase system approaching the triple point. At high under-cooling, the free energy presents two minima of equal height corresponding to two phase coexistence. Moving along the sublimation line the system remains at two phase coexistence, but a metastable minimum gradually emerges as under-cooling decreases. Eventually, the three minima adopt the same free energy at a triple point. (b) The corresponding order parameter profiles, showing the gradual enrichment of the metastable phase as under-cooling decreases. Lines from violet to red show the system behavior in order of decreasing under-cooling, as indicated by the black arrows.
@@ IV Wetting physics
where $\Pi=-dg(h)/dh$ is the disjoining pressure as introduced by the Russian school of colloidal science.Derjaguin (1987); Churaev (1988); Henderson (2005) In the geophysical community, $\Pi(h)$ is often known as the thermo-molecular force.Wettlaufer (1999a); Dash, Rempel, and Wettlaufer (2006)
This result shows that the film thickness may be considered as depending on the single variable $\Delta p$, which characterizes the approach of the vapor to liquid-vapor coexistence. Since the vapor at pressures close to the sublimation line behaves virtually as an ideal gas, $\Delta p(T)$ can be accurately approximated to $\Delta p(T)=\rho_{l}k_{B}T\ln p_{{sv}}(T)/p_{{lv}}(T)$, where $\rho_{l}$ is the density of the bulk liquid, $k_{B}$ is Boltzmann’s constant and $p_{\alpha v}$ is the coexistence vapor pressure over $\alpha={s,l}$.Llombart et al. (2020); Sibley et al. (2021) This shows explicitly that the equilibrium premelting thickness need not be measured strictly in terms of a distance away from the triple point. Rather, it is dictated by the distance of the sublimation line to the metastable prolongation of the condensation line. In water, these two lines run very close together all the way down to 0 K, an observation that could explain the unexpected report of ice premelting at very low temperature.Lin et al. (2023) Assuming Clausius-Clapeyron behavior for the vapor pressures, this equation simplifies to $\Delta p=\rho_{l}\Delta h_{{sl}}(T-T_{t})/T_{t}$, where $\Delta h_{{sl}}$ is the enthalpy of melting. The result is quantitatively correct down to $10^{\circ}$ C, whereupon it starts flattening but remains qualitatively correct for at least $50^{\circ}$ C.Baran et al. (2024)
@@
The reader might wonder what is the point of mapping $h_{e}$ to $\Pi(h)$ via $\Delta p$, if the experimental measurements provide right away $h_{e}$ as a function of temperature, which is a simpler variable. The answer is that, once $\Pi(h)$ is known, Eq.(8) shows that $h_{e}$ is a natural function of the single varible $\Delta p(T,p)$. This allows to predict equilibrium thickness for arbitrary thermodynamic states from knowledge of a set of measured states. For the case of surface premelting, knowledge of $\Pi(h)$ obtained for equilibrium states along the sublimation line, allows to predict the properties of premelting films in moderately out of equilibrium conditions of arbitrary temperature and pressure, as we shall discuss later in Section IX.Sibley et al. (2021)
@@
Figure 6: Characterization of film structure. Interfacially premelted films in widely different conditions but equal disjoining pressure are very similar. Profiles shown in dashed lines were obtained at $T=266$ K and $p=1$ bar; The ones in a continuum line are obtained at $T=262$ K and $p=400$ bar. The region of density profiles with a majority of liquid like molecules is depicted in either orange or blue; that where the majority is formed by solid like molecules is shown in green and black. Reproduced with permission from Baran et al.Baran, Llombart, and MacDowell (2025) J. Phys. Chem. C 129 4614 (2025). Copyright ©2025 American Chemical Society.

### v2

@@ I Introduction
On the 12th of March 1936 a team of Japanese researchers led by Ukichiro Nakaya succeeded in creating the first ever artificial snow crystal in a snow chamber under controlled conditions.Nakaya (1954) After several years of further research, the results were summarized and collected in the diagram that now bears his name.Nakaya (1954); Kobayashi (1957); Frank (1982); Furukawa and Wettlaufer (2007); Libbrecht (2022) The Nakaya diagram maps crystal habits as a function of temperature and humidity of the surrounding atmosphere (Figure 1). The complex structures that form as supersaturation is increased can be relatively well understood under the framework of crystal growth science.Frank (1982); Nelson (2001); Libbrecht (2022) However, the surprising transformation of solid hexagonal prisms from plate like to columnar like and back to plate like as temperature decreases has remained a mystery ever since.Frank (1982); Furukawa and Wettlaufer (2007); Libbrecht (2022)
@@
Kuroda went on to conduct ellipsometry and x-ray diffraction studies that were among the first few direct evidences for the presence of a quasi-liquid layer of melt water on the ice surface.Furukawa, Yamamoto, and Kuroda (1987); Kouchi, Furukawa, and Kuroda (1987) Following these studies, the field experienced a surge of renewed activity, with numerous measurements and estimates of the quasi-liquid layer thickness carried out using a variety of modern spectroscopic and mechanical techniques.Kouchi, Furukawa, and Kuroda (1987); Elbaum, Lipson, and Dash (1993); Lied, Dosch, and Bilgram (1994); Dosch, Lied, and Bilgram (1995, 1996); Wei, Miranda, and Shen (2001); Döppenschmidt and Butt (2000); Bluhm et al. (2002); Sadtchenko and Ewing (2002); Gelman Constantin et al. (2018); Mitsui and Aoki (2019)
@@
Interestingly, once the presence of liquid water droplets on the ice surface was confirmed by direct optical microscopy imaging, Faraday’s original hypothesis came under renewed scrutiny.Knight (1996) More recently, Sazaki and collaborators–at the Institute for Low Temperature Science, founded by Nakaya and later home to Kuroda–have raised serious concerns about this hypothesis,Sazaki et al. (2012); Asakawa et al. (2016); Murata et al. (2016); Sazaki et al. (2022), while studies on the ice surface using macroscopic probes have lead Bonn et al. to question the significance of ice premelting altogether.Liefferink et al. (2021); Demmenie et al. (2022, 2025)
@@
In this perspective article I discuss how a selected blend of concepts from these wide range of fields together with the help of computer simulations,Benet et al. (2016, 2019); MacDowell (2019); Llombart et al. (2019, 2020); Llombart, Noya, and MacDowell (2020); Sibley et al. (2021); Luengo-Márquez and MacDowell (2021); Luengo-Márquez, Izquierdo-Ruiz, and MacDowell (2022); Baran et al. (2022); Baran and MacDowell (2024); Baran et al. (2024); Baran, Llombart, and MacDowell (2025) allows to interpret and reconcile a large body of apparently conflicting empirical evidence into a unified and consistent picture of the structure and dynamics of premelting films. This includes some of the most relevant puzzles reported in recent reviews:Slater and Michaelides (2019); Nagata et al. (2019); Sazaki et al. (2022) How can the observation of elementary steps be reconciled with the presence of significant premelting? why do liquid droplets appear well above water-vapor saturation? And most notably, what is the reason for the alternation of crystal habits in the Nakaya diagram?
Often, this requires delving into rather intricate and subtle discussions. The readers less interested in the finer details of wetting physics and crystal growth can get an overall impression of the field by going through sections II, VI, IX, XI and XII only, while a read through section XII may be sufficient to get a broad picture of the current understanding and future challenges.
Figure 1: Diagram of snow crystal morphology, a descendant of the early Nakaya diagram due to Kobayashi.Kobayashi (1957) The figure displays the habit of ice crystals grown in the atmosphere as a function of temperature and excess water vapor density over ice saturation. The origin of the complex shapes that appear as saturation increases are relatively well understood. The reason why simple hexagonal prisms grown at low vapor saturation change from plates, to columns and back again to plates as temperature decreases has remained a mystery ever since Nakaya first published his results. Reproduced from Furukawa and Wettlaufer, Phys. Today 60 70 (2007),Furukawa and Wettlaufer (2007) with the permission of AIP publishing .
@@ II Modern empirical evidence of ice premelting
The experimental measurement of ice premelting thickness has been reviewed in a large number of papers.Elbaum, Lipson, and Dash (1993); Petrenko (1994); Sadtchenko and Ewing (2002); Bartels-Rausch et al. (2014); Asakawa et al. (2016); Gelman Constantin et al. (2018); Slater and Michaelides (2019) All of them agree that the data exhibit enormous dispersion, with discrepancies spanning several orders of magnitude. However, the large extent of scatter can be reduced significantly by a judicious selection of experimental results. As noted in Ref.Dash, Rempel, and Wettlaufer (2006) there is no reason to expect that the premelting film thickness should be equal for surface and interfacial premelting. Indeed, computer simulations show that interfacial premelting structure is similar to that of surface premelted films only for very hydrophobic substrates.Baran, Llombart, and MacDowell (2025) Considering the case of surface premelting alone reduces significantly the number of available measurements.Golecki and Jaccard (1978); Beaglehole and Nason (1980); Furukawa, Yamamoto, and Kuroda (1987); Elbaum, Lipson, and Dash (1993); Dosch, Lied, and Bilgram (1995); Döppenschmidt and Butt (2000); Sadtchenko and Ewing (2002); Gelman Constantin et al. (2018); Mitsui and Aoki (2019) However, out of all these studies, the early results by Golecki and Jacard using proton backscattering are clearly outliers,Golecki and Jaccard (1978) and can be ruled out. Furthermore, studies using Atomic Force Microscopy are necessarily invasive, and are prone to capillary condensation in between the tip and the surface,Pickering et al. (2018) so they are left out too. A final selection accounting for all the remaining measurements since 1990, using widely different non-invasive techniques such as ellipsometry, x-ray scattering, and infra-red, sum frequency or fluctuation spectroscopy lead to a rather different picture, with results differing by at most one order of magnitude for temperatures below about $1^{\circ}$C (Figure 2). The scatter remains bracketed within one order of magnitude even up to temperatures of $0.1^{\circ}$C if the results of Ref.Dosch, Lied, and Bilgram (1996) are discarded. All of these experimental measurements confirm that a premelting layer appears on the ice surface, with thicknesses that increase from the sub-nanometer scale to at most 10 nm in the temperature interval between 20-20 and $1^{\circ}$C. This appears somewhat higher but within reasonable agreement with computer simulation results for different water models, which consistently show the formation of a disordered liquid layer on the ice surface.Conde, Vega, and Patrykiejew (2008); Pfalzgraff, Neshyba, and Roeselova (2011); Limmer and Chandler (2014); Benet et al. (2016); Sánchez et al. (2017); Louden and Gezelter (2018); Qiu and Molinero (2018); Pickering et al. (2018); Kling, Kling, and Donadio (2018); Benet et al. (2019); Llombart et al. (2019, 2020); Llombart, Noya, and MacDowell (2020) Indeed, the extent of ice premelting can be quantified in computer simulations using suitable order parameters that allow to distinguish whether a water molecule is found in a liquid like or solid like environment (Figure 3).Lechner and Dellago (2008); Nguyen and Molinero (2015) Using such parameters, simulation studies report film thicknesses that lie in the sub-nanometer range up to about $1^{\circ}$C below the model’s triple point, and then grow somewhat above the nanometer thickness at higher temperatures.Conde, Vega, and Patrykiejew (2008); Limmer and Chandler (2014); Qiu and Molinero (2018); Pickering et al. (2018); Llombart et al. (2019, 2020); Llombart, Noya, and MacDowell (2020); Berrens, Bononi, and Donadio (2022)
Hower, notice that in strict sense the direct comparison between computer simulations and experiments can only be done for studies probing uncontaminated and atomically smooth ice surfaces with well defined orientation. This requires dedicated preparation of single mono-crystals grown by layer by layer deposition from the vapor.Furukawa, Yamamoto, and Kuroda (1987); Elbaum, Lipson, and Dash (1993); Sazaki et al. (2012); Asakawa et al. (2015); Murata et al. (2016); Libbrecht (2022) In most laboratories, measurements have been performed from rapid vapor deposition,Bluhm et al. (2002); Gelman Constantin et al. (2018) or water freezing,Beaglehole and Nason (1980); Sadtchenko and Ewing (2002); Mitsui and Aoki (2019) which produces polycrystalline samples with uncontrolled smoothness unless extra care is exercised.Dosch, Lied, and Bilgram (1995) In practice, it seems from Figure 2 that the premelting film thickness remains constrained within one order of magnitude irrespective of the smoothness or surface orientation. Particularly, the current uncertainties in experiments do not seem to settle whether the film thickness is significantly different in basal or prism surfaces, but results from computer simulations suggest this difference is constrained within tenths of a nanometer.Conde, Vega, and Patrykiejew (2008); Llombart et al. (2020); Llombart, Noya, and MacDowell (2020); Baran et al. (2024)
@@ III A toy model of (complete) surface melting
Figure 4: Phase diagram of water in the neighborhood of the triple point. Full black lines illustrate the equilibrium phase boundaries, including the sublimation line separating ice from vapor and the condensation line, separating liquid water from vapor. The black dashed line is the metastable prolongation of the condensation line. Equilibrium premelting occurs for a path along the sublimation line, as indicated by the blue arrow. The equilibrium thickness of the premelting film is dictated by the distance of the sublimation line to the metastable prolongation of the condensation line, as indicated by the green segment. The red line indicates a non-equilibrium path, with three different regimes as described in Section IX. In region a, below the kinetic condensation line, shown in dashed blue, ice grows but the premelting film remains in a steady state of constant thickness. In region b, above the kinetic condensation line but below the kinetic spinodal line shown in dashed red, the premelting film remains in steady state but vapor can condense and form droplets atop. In region c, above the kinetic spinodal line, condensation occurs faster than freezing, the premelting film diverges and ice freezes in a wet mode.
@@
Premelting, viewed as an equilibrium phenomenon, is the study of how the metastable liquid phase ($l$) intrudes at the interface between the coexisting bulk solid ($s$) and vapor phases ($v$) in the vicinity of a metastable liquid-vapor coexistence line (Fig. 3) . For many substances, this can only occur as the system approaches the triple point. For the special case of ice, however, notice the metastable prolongation of the liquid-vapor coexistence line runs very close to the sublimation line all the way from the triple point down to very low temperatures .
@@
Within the square gradient approximation, the excess surface free energy of the interface , $\Delta\omega$, may be obtained in terms of $\phi(z)$, as a sum of bulk free energies adopted by the system along the interface. An additional square gradient term accounts for the cost of creating the inhomogeneity:
@@
A simple way to remedy this deficiency is to modify the bi-quadratic potential to account for the intrusion of a third phase with an order parameter in between those of the bulk solid and liquid phases, i.e. $f(\phi)=(1-\phi^{2})^{2}(t+\phi^{2})$, where $t\propto 1-T/T_{t}$ is a measure of the under-cooling (Figure 5-a) . The Euler-Lagrange equation, Eq.(2) can be solved exactly for this potential, yielding:Lipowsky and Speth (1983)
@@
This result nicely describes the essence of premelting. For deep under-cooling, $t$ is much larger than $1-\tanh^{2}(\kappa z)$ for all $z$, and Eq.(4) becomes virtually equal to the sharp $\tanh(\kappa z)$ interfacial profile that is typical for two phases at coexistence. When the system approaches the triple point, however, $t\to 0$, the order parameter profile adopts a value $\phi(z)\approx\sqrt{t}$ close to zero as long as $1-\tanh^{2}(\kappa z)$ in the denominator remains larger than $t$. This occurs in an increasingly larger range of $z$ values as $t$ vanishes. Accordingly, the model predicts a gradual increase of the liquid phase in between the solid and the vapor phases on approaching the triple point (See Figure 5 -b ).
@@
Using the above results it is also possible to discuss what precisely can be understood as the onset of premelting.Qiu and Molinero (2018); Slater and Michaelides (2019); Nagata et al. (2019) The model shows that premelting appears in a continuous fashion, as a gradual flattening of the interfacial profile at the interface position $z=0$, i.e., there is no particular thermodynamic signature in the density profile of the solid-vapor interface signaling a premelting threshold. The best that can be done is to notice that the flattening of the interfacial density is the result of the stabilization of the liquid phase at a metastable minimum of the free energy surface. This is the only feature that can be identified as the theoretical onset of surface premelting. However, the signature of the metastable minimum is hardly detectable in the order parameter profile until the minima of $f(\phi)$ become sufficiently close in energy, i.e. $t\approx 0$. Therefore, for practical matters the onset cannot be identified until the change in slope and the emerging liquid film at $z=0$ is above the chosen experimental probe’s detection threshold.
[…]
Figure 5: Toy model of surface premelting. (a) Free energy of a two-phase system approaching the triple point. At high under-cooling, the free energy exhibits two minima of equal height, corresponding to coexistence of the two phases. Along the sublimation line, a metastable minimum gradually emerges as under-cooling decreases. At the triple point, all three minima reach the same free energy. (b) Corresponding order parameter profiles, showing the gradual enrichment of the metastable phase as under-cooling decreases. Lines from violet to red indicate system behavior in order of decreasing under-cooling, as highlighted by the black arrows.
@@ IV Wetting physics
where $\Pi=-dg(h)/dh$ is the disjoining pressure as introduced by the Russian school of colloidal science.Deryagin (1987); Churaev (1988); Henderson (2005) In the geophysics community, $\Pi(h)$ is often known as the thermo-molecular force.Wettlaufer (1999a); Dash, Rempel, and Wettlaufer (2006)
This result shows that the film thickness may be considered as depending on the single variable $\Delta p$, which characterizes the approach of the vapor to liquid-vapor coexistence. Since the vapor at pressures close to the sublimation line behaves virtually as an ideal gas, $\Delta p(T)$ can be accurately approximated to $\Delta p(T)=\rho_{l}k_{B}T\ln p_{{sv}}(T)/p_{{lv}}(T)$, where $\rho_{l}$ is the density of the bulk liquid, $k_{B}$ is Boltzmann’s constant and $p_{\alpha v}$ is the coexistence vapor pressure over $\alpha={s,l}$.Llombart et al. (2020); Sibley et al. (2021) This shows explicitly that the equilibrium premelting thickness need not be measured strictly in terms of a distance away from the triple point. Rather, it is dictated by the distance of the sublimation line to the metastable prolongation of the condensation line. In water, these two lines run very close together all the way down to 0 K, an observation that could explain the unexpected report of ice premelting at very low temperature.Lin et al. (2023) Assuming Clausius-Clapeyron behavior for the vapor pressures, this equation simplifies to $\Delta p=\rho_{l}\Delta h_{{sl}}(T-T_{t})/T_{t}$, where $\Delta h_{{sl}}$ is the enthalpy of melting. The result is quantitatively correct down to $10^{\circ}$ C, but remains qualitatively correct for at least $50^{\circ}$ C.Baran et al. (2024)
@@
The reader might wonder what is the point of mapping $h_{e}$ to $\Pi(h)$ via $\Delta p$, if the experimental measurements provide right away $h_{e}$ as a function of temperature, which is a simpler variable. The answer is that, once $\Pi(h)$ is known, Eq.(8) shows that $h_{e}$ is a natural function of the single variable $\Delta p(T,p)$. This allows to predict equilibrium thickness for arbitrary thermodynamic states from knowledge of a set of measured states. For the case of surface premelting, knowledge of $\Pi(h)$ obtained for equilibrium states along the sublimation line, allows to predict the properties of premelting films in moderately out of equilibrium conditions of arbitrary temperature and pressure, as we shall discuss later in Section IX.Sibley et al. (2021)
@@
Figure 6: Characterization of film structure on substrates of different wettability as indicated by the contact angle, $\theta$. Interfacially premelted films in widely different conditions but equal disjoining pressure are very similar. The density profiles shown in dashed lines were obtained at $T=266$ K and $p=1$ bar; The ones in a continuum line are obtained at $T=262$ K and $p=400$ bar. The region of density profiles with a majority of liquid like molecules is depicted in either orange or blue; that where the majority is formed by solid like molecules is shown in green and black. Reproduced with permission from Baran et al.Baran, Llombart, and MacDowell (2025) J. Phys. Chem. C 129 4614 (2025). Copyright ©2025 American Chemical Society.